OpenStudy (hrazor):

The areas of two similar tetrahedrons are 24 sq. in. and 96 sq. in. Find the ratio of their volumes.

Directrix (directrix):

If two similar tetrahedrons have areas in the ratio of 24 to 96, then the scale factor of the two tetrahedrons is sqrt(24/96) =sqrt(1/4) = 1/2.

Directrix (directrix):

The cube of the scale factor is the ratio of the volumes.

Directrix (directrix):

(1/2)^3 is the ratio of the volumes.

OpenStudy (hrazor):

Thanks.